- 3/27-4/10: Fixing PR2
- 4/4 -4/5: Gluing
- 4/12: Closing and pump down
3/19: Cooling stop
3/25: Venting start
3/27: Opning chamber
3/27~: Fixing thermal touch
4/2~: closing chamber and start pumping down.
- 4/12: Cooling starts.
3/25: closing and pump down
- 4/11: Cooling start
3/26: Venting start
3/28~: Fixing short inside chamber
4/3: Fixing F0 actuator
4/4~: closing chamber and start pumping down.
- 4/9: Cooling starts.
- 4/8~ or 4/15~ : Night work starts, daily meeting will be moved to 1PM
- Summary of work will last for a while in the morning meeting.
- April: Measurement of TF, payload diagonalization on ITMX
- End of April~: DRMI commissioning
- End of April~: Y arm commissioning
- ~End of May: FPMI commissioning
- Official night work starts 4/8 or 4/15.
- Saturday and Sunday will be on laser hazard all the area.
- Sat., Sun.: iMC alignment (Kokeyama)
- Sat.: ETMX (Lucia)
- Sat.: ETMY (Suzuki)
Report of work
- Type Bp: Total 3 flags were glued on PR2 and PRM. Jigs will be released today.
- CRY: It was found that on of the PD sensor at ETMX cryostat has no output. Now Ushiba-kun is checking it.
- PEM, VAC: Vacuum values of center area have been added to monitor by Yokozawa-kun.
- VIS: All VIS will be finally checked today and next week.